OpenStudy (anonymous):

If John mows the lawn in 10 hours, Tim mows the lawn in 8 and together they complete the lawn in 4 8/18 hours. How much lawn would be covered in 1 hour by John, by Tim, and together?

OpenStudy (anonymous):

Let x be the faction of a whole lawn that can be mowed in one hour. Solve the following equation for x:\[\frac{1}{4+\frac{8}{18}}=\frac{x}{1}\]The expression above can be verbalized as follows: "If the lawn can be mowed by John and Tim in 4 and 8/18 ths hours, then they can mow x of it in one hour.\[x=\frac{9}{40} \text{ of one lawn} \]

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

You are welcome.